AnimeGhoul8863:

Consider the function f(x) = x2 − 2x + 6. Find the value of f(x) when x = 3 and select the correct answer below. 1 3 6 9

AnimeGhoul8863:

@TheSmartOne

NarWhalGod:

It's obviously unicorns duhhhh

BenLindquist:

f(x) = x^2 - 2x + 6

AnimeGhoul8863:

^yes

NarWhalGod:

Which one

BenLindquist:

f(3) = 3^2 - 2 x 3 + 6 = 9 - 6 + 6 = 9

TheSmartOne:

This question is really simple, they said x=3, and all you have to do is plug it in to find out what f(3) is